what is the solution of x+y=4 and 2x-y=-5
Found 2 solutions by longjonsilver, Abbey:
Answer by longjonsilver(2297)

Go to the homework section on this site and look at the lessons i have done on linear equations. It explains them.

add the 2 equations together here and you get 3x = -1, so x=-1/3. put this in either equation to find y. THEN put both answers in the other equation to check your answer is correct!

jon

Answer by Abbey(339)

This problem is perfect for the substitution method:
Rewrite x + y = 4 as
y = 4-x by subtracting x from both sides of the equation.
then use this (4-x) in the second equation for y:
2x-y=-5 as
2x-(4-x)=-5
Clear the parentheses:
2x-4+x=-5
Combine like terms:
3x-4=-5
3x-4+4=-5+4
3x=-1
Divide both sides by 3 to isolate the variable:
x=-1/3
By putting the x value back into either equation:
-1/3+y=4
y=4 2/3
After putting both x and y into the equations to verify the results:
-1/3 + 4 1/3 = 4
2(-1/3)-4 1/3 = -5
The solution set is:
(-1/3,4 1/3)